The heat produced in a given resistance in a time by  the sinusoidal current I°sinwt will be the same as that of a steady current of magnitude nearly

Grade:12th pass

1 Answers

Susmita
5 years ago
The heat produced is
H=I2Rt .
For the sinusoidal current within a time T=2pi/w
H=\int_{0}^{T}Io^2sin^2\omega t R dt
=\frac{Io^2 RT}{2}
So the steady current magnitude is
 \frac{Io}{\sqrt{}2}